OpenStudy (keythemusician):

~~Fan + Medal~~ Can someone please explain how a four-term polynomial is factored by grouping and when a quadratic trinomial can be factored using this method. Include examples in your explanation.

OpenStudy (3mar):

For a four-term polynomial factorization: \[x^3+2x^2-x-2=0\] \[(x^3-x)+(2x^2-2)=0\]this is the step of grouping, then: take common factors: \[x(x^2-1)+2(x^2-1)=0\] take \[(x^2-1)\] as a common; \[(x^2-1)(x+2)=0\] then expand \[(x^2-1)\]: \[(x-1)(x+1)(x+2)=0\]

OpenStudy (keythemusician):

Thank you!

OpenStudy (jiteshmeghwal9):

What if (x^2-1) ain't a common factor

OpenStudy (3mar):

Excuse me, what is "quadratic trinomial can be factored using this method"?

OpenStudy (keythemusician):

Um I'm not sure

OpenStudy (3mar):

"What if (x^2-1) ain't a common factor" you search for the best composition in order not to be in that situation. If you find that there is no common, then you can complete a quadratic ,,,,, another way

OpenStudy (3mar):

"Um I'm not sure" how can I solve such thing you don't sure about ;;)

OpenStudy (keythemusician):

It just what the question asked, I'm searching in the lesson for the answer

OpenStudy (keythemusician):

I figured it out a quadratic trinomial is just x^2 + bx + c and they want it to be solved by factoring by grouping.

OpenStudy (3mar):

HOW "solved by factoring by grouping."?

OpenStudy (keythemusician):

1. Check for GCF 2. Split the middle term 3. Factor by grouping 4. Check by distributing